Two identically charged one-peso coins are 1.5 m apart on a table. What is the charge of the coins if each of them experiences a repulsive force of 2.0 N?


1
Expert's answer
2022-03-03T12:05:50-0500

The Coulomb's law says

"F=k\frac{q_1q_2}{r^2}=k\frac{q^2}{r^2}"

The absolute value of charge

"q=\sqrt{Fr^2/k}\\
q=\sqrt{\frac{2.0*1.5^2}{9*10^9}}=2.2*10^{-5}\:\rm C"